OpenStudy (anonymous):

The position (in feet) of an object is given by s(t) = (25- t 2) 3 where t ~ 0 is the time elapsed in seconds. Find the time at which the acceleration of the object is zero. When is the object at rest?

OpenStudy (anonymous):

please help!

OpenStudy (sriram):

acc=0 means the double derivative=0 v=3(25-t^2)^2 (-2t) a=3*2(25-t^2) (-2t) * 3(25-t^2)^2 (-2)=0 solve it

OpenStudy (sriram):

i now feel expanding it and solving it would have been easier

OpenStudy (anonymous):

what does it mean when it says when the object is at rest?

OpenStudy (sriram):

the object it at rest means the velocity =0 for that put the v we got up =0

OpenStudy (sriram):

v=0 gives 25-t^2=0 hence t=5

OpenStudy (anonymous):

gotcha!! thank you!!!!

OpenStudy (sriram):

i guess we get the same result for acc.

OpenStudy (anonymous):

i'm still not getting it for some reason....for v(t) i have -6t(25-t^2)^2 is that correct?

OpenStudy (sriram):

yes

OpenStudy (anonymous):

then you set that to 0? but how do you solve for t?

OpenStudy (sriram):

look now -6t(25-t^2)^2=0 hence for it to be 0 either t=0 or 25-t^2=0 accordin to ur Question t is not 0 hence the second case

OpenStudy (anonymous):

ohhh okay that makes sense now!!

OpenStudy (anonymous):

and same for a(t)=24t(25-t^2) either t=0 or +/- 5? so you don't distribute 24t to what is inside the parenthesis you just set both to zero?